Home‎ > ‎Crew 143‎ > ‎


Engineering Report


Ian Siversides


In the middle of the night, a second rat found the trap and is now disposed of.


Trailer water tank: 35/40 in

Propane tank: 16%

Generator diesel: 55%


Space pack 1 had a blown 5 A fuse. This has been rectified.

Space pack 3’s power drains quickly.

Space pack 5 still in pieces (missing battery).

All other space packs functioning nominally (5 serviceable space packs).


The defogging device consisting of magnets wrapped in fabric are no longer necessary as the visors are prepared with a film of soap, and silica packets are placed in the helmets.


Measurements for the 7th crew’s quarters were passed on to Crew 144.


It is windy today and the Hab is noticeably cooler. An adjustment to the thermostat was not found to be necessary. It is currently a comfortable 20C.


The power was noticed to flicker > 10 times this afternoon. Presumably gusts of wind are the cause.


The water heater needed a restart - it is now functioning normally.


Yesterday’s waster use: 20 L/person (1 sponge bath).

Today’s current waster use (will increase): 36 L/person (1 shower).


Even with our use of extreme caution, the toiled backed up for the first time this afternoon. Our water consumption after lunch was 12 L/person; we could have set another record today.